OpenStudy (anonymous):

What is the equation of the circle with center at (-3, 0) and diameter 20?

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

So I gave you the stuff to find out last time when I told you the equation of a circle. Show me your best guess brony.

OpenStudy (anonymous):

this is different! whats the difference between h and k, and x and y? how do i figure this out? at least give me the equation.

OpenStudy (kainui):

(x-h)^2+(y-k)^2=r^2 Is the equation.

OpenStudy (kainui):

if (h,k) is the center of the circle and r is the radius, then you should be able to plug in the information you have to figure it out, give it a try.

OpenStudy (anonymous):

no, the equation with the numbers inserted. cuz i dont know if the coordinates are for x an y, or h and k?

OpenStudy (kainui):

(x-h)^2+(y-k)^2=r^2 Is the equation. (h,k) is the coordinates of the center of the circle. r is the radius You are given: circle with center at (-3, 0) and diameter 20 This is all the information you need to look at, see if you can at least make a guess and try, it doesn't matter if you mess up, I'm here to help.

OpenStudy (anonymous):

(x--3)^2+(y-0)^2=r^2

OpenStudy (anonymous):

(x + 3)2 + (y - 0)2 = 400

OpenStudy (anonymous):

Diameter is 29therefore radius is 20/2=10

OpenStudy (kainui):

Almost, remember they give you the diameter, but the equation has radius! Other than that, it looks fine. I'm not sure, but it looks like you dropped your exponents too.

OpenStudy (anonymous):

20* sorry about that